Suppose the cost function is C(Q) = 50 + Q − 10Q2 + 2Q3. At 10 units of output, the average cost curve is


a. at the maximum level.


b. in the declining stage.


c. in the increasing stage.


d. at the minimum level


1
Expert's answer
2021-11-18T10:24:38-0500

TC = 50 + Q - 10Q2 + 2Q3

Average Total Cost Function

"ATC = \frac{TC} { Q}"

"ATC = \frac{(50 + Q - 10Q2 + 2Q3)}{ Q}"





So from the above table, from 1 to 4 units the average total cost is decreasing and from 5 units onwards the average total cost is increasing.

Hence at 10 units of output, the average total cost curve is in the increasing stage



The correct answer is c. in the increasing stage.